Explain powar arpresston in phasor domain VA = V z real componant +rectve componant
Answer #1

Given the question is to explain the power expression, I assume the main doubt would be the explanation behind using a conjugate for calculating power.

Consider a case where we are not using the conjugate and power = V*I

elaborating it, Power = V\angle \alpha I \angle \beta and this product turns out to be VI (cos(\alpha + \beta) + j cos( \alpha - \beta)). For suppose if \alpha + \beta > 90o then power comes out to be of the form -P + jQ (P and Q are just representations, they hold no significant values)which is unacceptable for the real part to be a negative quantity

On the other hand, if we use the conjugate of current which is I*, then Power = V \angle \alpha I- B and this turns out to be VI (cos (\alpha -  \beta) - j cos (\alpha + \beta)). Here, we have to analyse the problem by considering to situations,

Case 1: if \alpha > \beta(Voltage leads current), then the difference is positive and less than 900 hence the Power comes out to be of the for P - jQ which by the fundamentals we know is an inductive one.

Case 2: if \alpha < \beta(Voltage lags current). Could you guess what would happen? Yes, it becomes a capacitative one.

Hence this summarizes the use of phasor in the power expression and precisely why we use a conjugate for our calculation of power.